OpenStudy (anonymous):

What is the sale price of a cell phone that is on sale for 15% off if the regular price is $75? $60.00 $63.75 $64.25 what equation should i use to figure this out

OpenStudy (anonymous):

Start out with the equation (15/100) = (x/75) Cross multiply and you get: 100x = 1125 Simplify for x: x = 11.25 So now you have $11.25 off. Subtract that from the original price: 75-11.25 = 63.75 Your answer is 63.75

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

yessir! Hope you understand it

OpenStudy (anonymous):

now that u explained it yes

OpenStudy (anonymous):

cool!